OpenStudy (anonymous):

Simplify the expression -- square root of 81x^6y^-4

OpenStudy (anonymous):

\[\sqrt{81x ^{6}y ^{-4}}\]

OpenStudy (angela210793):

=9(X)^3y^(-2)

OpenStudy (angela210793):

=9(X)^3*y^(-2)